### v2

# Thermo-mechanic invariant in anisotropic plasma with heat fluxes

Abstract
We investigate linear perturbations in an incompressible, weakly collisional, anisotropic plasma, taking heat fluxes into account. Our analysis identifies a new class of time-invariant perturbations that encompass heat flux variations and aperiodic longitudinal perturbations in the magnetic field and flow velocity. Under the Chew-Goldberger-Low (CGL) approximation, which neglects low-frequency heat flux effects, these perturbations are purely thermal and decoupled from magnetic and velocity fields. When heat fluxes are included, these perturbations gain kinematic and magnetic components, leading to the identification of a “thermo-mechanic mode.”
The nature of aperiodic linear modes suggests that they may reach higher amplitudes in perturbation structures, that can be revealed as filamentary formations. We discuss the potential implications of these filaments on the large-scale dynamics of flows and on the nonlinear energy balance within small-scale MHD turbulence. Our findings indicate that thermo-mechanic invariant may play a significant role in modulating turbulence energy transfer dynamics, potentially impacting both local and global flow behavior in weakly collisional anisotropic plasmas.
I Introduction
A weakly collisional, rarefied plasma may exhibit anisotropic behavior under the influence of the Lorentz force resulting from an external magnetic field. In ionized media, the particle gyration radius around the magnetic field lines can be shorter than the mean free path of particle collisions, resulting in directional anisotropy: the effective pressure and temperature of the plasma differ along and perpendicular to the direction of the magnetic field. The study of weakly collisional anisotropic plasmas is crucial for understanding the behavior of various astrophysical and laboratory plasma environments, where the interplay between effects of magnetic field and heat fluxes influences both stability and energy distribution. Such flows are often found in laboratory plasma experiments as well as solar wind, stellar and galactic outflows, and various large-scale ionized flows throughout the Universe.
A comprehensive mathematical description of anisotropic plasma flows requires a proper Boltzmann-Vlasov kinetic description. On the other hand, fluid models can offer significant simplifications while still capturing essential large-scale phenomena. Building on the success of Magnetohydrodynamic (MHD) theory in modeling plasma flows within a fluid approximation, Chew, Goldberger, and Low (CGL) developed an extended framework for anisotropic flows[1, 2]. CGL approximation provides a fluid-based model for anisotropic plasmas by introducing a double adiabatic closure for pressures parallel and perpendicular to the magnetic field direction. Although successful in describing many large-scale plasma flows, the CGL approximation neglecting a so-called “pressure transport” terms neglects key low-frequency heat flux effects, leading to limitations in the study of phenomena such as flow stability and thermal transport.
Addressing limitations of CGL approximation requires further development of the closure model, when heat flux variables are explicitly incorporated into a higher-order equation of state. [3, 4, 5, 6, 7]. The derived system is commonly referred to as a closure model with heat fluxes. This model significantly enhances the fluid description of weakly collisional plasmas in the low-frequency limit, where thermal effects are particularly pronounced. Consequently, the heat flux fluid model is particularly well-suited for characterizing low-frequency and aperiodic phenomena in such media. Indeed, several notable advancements have been achieved in our understanding of various physical phenomena in collisionless plasmas using the proposed approximation[8, 9, 10, 11, 12, 13, 14, 15]. These advancements underscore effectiveness of the proposed model in capturing intricate behaviors of plasma, such as wave propagation, turbulence characteristics, and energy transfer mechanisms. Furthermore, this approach can deepen our understanding of collisionless plasmas, ultimately contributing to more reliable predictions in astrophysical and laboratory plasma environments.
The present paper aims to study the anisotropic MHD fluid model with heat fluxes, specifically focusing on the linear dynamics of perturbations within the aperiodic limit. By employing a linear perturbation analysis in real-space coordinates, we investigate time-invariant conservative quantities and their associated perturbation fields. This methodology facilitates a comprehensive understanding of how heat fluxes influence the low-frequency dynamics of collisionless or weakly collisional magnetized plasma flows. Furthermore, we demonstrate that aperiodic linear modes can significantly affect the energy balance of turbulent flows, as well as the kinematic and magnetic characteristics of anisotropic plasma flows.
The paper is organized as follows: In Section II, we present the governing equations for a weakly collisional, anisotropic plasma, incorporating both pressure anisotropy and heat fluxes. We derive the linearized perturbation equations and analyze the conditions under which the thermo-mechanic invariant emerges. Section III summarizes the key findings of the paper and discusses potential applications of these results to the nonlinear global dynamics of plasma flows in various contexts.
II Physical Model
The dynamics of incompressible collisionless plasma with heat fluxes can be described within the anisotropic fluid approximation using the following set of equations [14, 15]:
that are complemented by the double adiabatic equations for the parallel () and perpendicular () to the magnetic field pressure components:
modified by the action of the parallel () and perpendicular () heat fluxes. Here is the unity vector in the direction of the magnetic field and
The close set of the fluid model can be obtained using the equations for heat fluxes as follows:
Here the parameter is employed to denote zero heat flux limit of the problem, effectively allowing for the reduction of the obtained results to the CGL limit:
The system of Equations (1-8) forms a closed set that governs the dynamics of weakly collisional plasma within the fluid closure model, accounting for the effects of heat fluxes. The apparent complication of the system as compared with the standard MHD fluid model comes from the fact that equation of state is now formed by Eqs. (5-8) introducing additional degrees of freedom.
II.1 Equilibrium state
For the equilibrium configuration of the plasma flow, we assume a uniform density () , constant flow aligned with the background uniform magnetic field (). For the simplicity of the description, we align the x-axis along the direction of the magnetic field:
and describe the plasma in the co-moving with the flow frame where . Background anisotropic pressure and heat flux parameters can be described in the components parallel and perpendicular to the magnetic field direction:
where plasma anisotropy parameter can be calculated as:
To parametrize the problem we may introduce the following well known physical parameters:
For the compactness of the notations we may introduce the normalized heat flux parameters:
parameters associated with the kinetic firehose mode:
and anisotropy induced difference between the perpendicular and parallel sound speeds:
It is important to note that both and can take a positive or negative values, depending on the microscopic stability conditions, such as the presence of the firehose instability[15, 16, 17, 18], or the nature of plasma anisotropy (e.g., when ).
II.2 Linear Perturbations
The dynamics of linear perturbations over the above described equilibrium state may be analyzed using the standard linearization procedure. For the shortness of the notations we introduce a generalized vector for the physical variables of the problem:
Hence, we may introduce generalized vectors representing equilibrium () and perturbed () components of the physical variables:
Having in mind the equilibrium state introduced by the physical model of our problem we derive:
Thus, by introducing linear perturbations over the background values and neglecting the nonlinear terms, we can derive a system of linear differential equations that describe the dynamics of perturbations in the anisotropic medium. To represent the system in the non-dimensional form we utilize the generalized vector with normalized physical variables
where
and introduce the typical length-scale of the problem and switch to the non-dimensional description using:
Hence, the dynamical equations can be represented in the non-dimensional matrix form as follows:
where is the 10th order matrix operator defined by:
where dots denote zero elements of the matrix. Using divergence free conditions for the perturbations of the velocity and magnetic field
we may derive stationary relation between parallel and perpendicular components of the pressure perturbations as follows:
where the following differential operators are used:
Hence, we may reduce the system by taking its derivative and substituting linear perturbations of , and using the above equations. Let us defined the reduced perturbation vector as:
Now, the temporal dynamics of linear perturbations may be described by:
where is the the 7th order matrix operator given by:
The non-stationary solutions for linear perturbation system of incompressible anisotropic flows is already analyzed using the 6 order dispersion equation[5]. Our eigenfunction analysis reveals that the 7th order matrix can describe an additional mode, which does not appear in the time variation analysis but can possess a non-zero, non-trivial solution.
II.3 Stationary Solution
The time invariant quantity of the system, if present, can be expressed as a linear combination of perturbations:
where is vector generator of the invariant quantity . The existence of a conservative quantity can be established through the presence of a nontrivial solution () of the following stationarity condition:
Using Eq. (27) in the stationarity condition (28) and assuming that generator is uniform both in space and time we may derive the following condition:
Hence, Eq. (25) leads to the realizability condition for the existence of the aperiodic mode:
Following the particular symmetry features of the matrix operator for our setting we may derive:
The existence of a general solution for the invariant in the form of for arbitrary perturbations requires:
Consequently, the 7th order system is further simplified to a 4th order problem, which possesses a non-trivial solution () for the vector generator components, expressed in the following form:
Here the operator is introduced for the shortness of the notations:
Hence, utilizing the vector generator derived from Eqs. (31-36), we can obtain the explicit form of the time-invariant variable as follows:
It appears that the perturbations of the linear stationary mode encompass heat fluxes in both parallel and perpendicular directions, as well as perturbations in velocity and magnetic field along the background magnetic field. Consequently, we can identify as the thermo-mechanic invariant of the linear system. Existence of such conserved quantity highlights the presence of stable, non-oscillatory modes, which may play a significant role in the redistribution of energy within weakly collisional anisotropic plasmas.
To trace this invariant in the CGL limit we can set and neglect the effects of heat fluxes by omitting the heat flux parameters and .Consequently, we obtain:
In this limit, it appears that possesses a purely thermal nature and does not involve the kinematic or magnetic characteristics of the plasma flow.
Based on the definitions provided above, we can express the thermo-mechanic invariant in the presence of heat fluxes () in a manner that highlights its magnetic and kinetic components as follows:
where
and
Equation (41) shows that an increase in the heat flux parameter enhances the magnetic field contribution to the invariant, which would otherwise be primarily dominated by the thermal () and parallel velocity () components. On the other hand, the invariant form reveals its inherently spatially compact nature. A uniform or nearly uniform perturbations, when
cannot excite the thermo-mechanic mode with a notable amplitude. As a result, perturbation field of the invariant can be characterized as localized filaments, which lead to the excitation of an aperiodic thermo-mechanic quantity that is conserved throughout the time evolution of the anisotropic flow.
To show an illustrative example of the thermo-mechanic invariant field we use 2D reduction to the plane. In this limit heat fluxes contribute into the thermal component of invariant as follows:
Assuming simplest heat flux perturbation field with power low spacial distribution contributing to the uniform in space invariant we may construct solution of the Eq. (48) as follows:
Figure 1 illustrates a specific configuration of the heat flux field of the thermo-mechanic invariant for a given set of parameters. The geometry of heat fluxes in the aperiodic mode appears to be strongly influenced by the anisotropy parameter .
III Summary
We study the linear perturbations in the anisotropic magnetised plasma flow with heat fluxes. Our findings suggest that such systems can sustain a new type of variable – time invariant perturbations that involve heat fluxes alongside with aperiodic longitudinal perturbations of both the magnetic field and flow velocity.
In the CGL approximation, this mode is purely thermal and decoupled from magnetic and velocity fields. However, when the heat fluxes are explicitly incorporated into the governing equations, we identify a new class of invariant perturbations with both kinematic and magnetic field components, which we refer to as the “thermo-mechanic mode.” This new linear mode reveals an important link between thermal and kinetic aspects of the plasma, indicating that the heat fluxes, often neglected in simpler models, plays a significant role in shaping the plasma dynamics, particularly in weakly collisional, anisotropic flows.
Our analysis shows that these thermo-mechanic modes persist as aperiodic, invariant quantities that may have profound implications for energy transfer and stability in weakly collisional plasma environments. These modes could influence the long-term behavior of turbulent plasma flows by redistributing energy between different degrees of freedom, including thermal, magnetic, and velocity fields.
We have derived the perturbation field of the thermo-mechanic mode in the coordinate space. It seems that the thermo-mechanic invariant is characterized by small-scale structures, such as filaments, within the perturbation field. This property may significantly alter the energy balance of turbulent flows and resulting turbulent spectrum in weakly collisional plasma flows.
In small-scale MHD turbulence, the governing physical processes exhibit significantly more diversity than those observed at larger scales[19, 20, 21, 22]. The linear conservative modes may add to this complexity: the mode effectively emerging at smaller scales suggests a shift in the nonlinear balance within the turbulence spectrum, indicating that the energy dynamics at these scales may be fundamentally altered. The MHD turbulence cascade is normally dominated by nonlinear couplings between different linear modes. These nonlinear interactions drive energy transfer between modes and a new aperiodic mode can give rise to a new energy exchange channels in the flow. At small scales, perturbations of the linear thermo-mechanic mode can potentially shift how energy is distributed across scales and redefine how turbulent energy dissipate. In this case, thermal effects can unexpectedly produce significant measurable impact, in addition to the classical dissipative effects. Thus, a new aspects of small-scale MHD turbulence dynamics that are essential for understanding energy transport and magnetic field evolution in astrophysical and laboratory plasma environments can be studied.
It remains to be determined whether the thermo-mechanic mode is purely a linear microscopic characteristic of perturbations or if it can also manifest at macroscopic scales. If the latter is true, it could lead to a direct connection between thermal perturbations and the longitudinal kinetic energy of velocity perturbations. This connection could substantially affect the macroscopic acceleration and deceleration of anisotropic magnetized plasma flows, a phenomenon that may have significant implications for the dynamics of solar and stellar wind flows.
In conclusion, our research adds an important aspect to a robust framework for investigating aperiodic thermal perturbations in anisotropic plasma. Future studies should aim to extend this model to explore nonlinear dynamics and assess how these perturbations affect large-scale plasma behavior and stability across various astrophysical conditions.
To definitively address these questions, assess the applicability of the acceleration mechanism in the context of solar and stellar winds, or analyze the thermal non-dissipative modifications of the small-scale turbulence spectrum in anisotropic flows, a nonlinear analysis is required. This can be achieved through direct numerical simulations of anisotropic flows with an appropriate configuration. Meanwhile, the invariant can be used as a benchmark value for verifying the accuracy of numerical codes during the analysis of the evolution of small amplitude waves.
Acknowledgements.
We acknowledge support from the Georgian National Astrophysical Observatory and note valuable discussions with Bidzina Shergelashvili.Data Availability Statement
The data that support the findings of this study are available within the article.